# Agentic Neurodivergence as a Contingent Solution to the AI Alignment Problem

### v1

@@ 6 Conclusions
In [17] the significant challenge of aligning AI to human values is highlighted, particularly when AI models are exposed to unethical content. They argue that simply filtering or restricting access to such content is not a sustainable solution. Instead, we need AI systems that can, like humans, understand and process diverse information, including unethical content, and still choose to act ethically. In fact, exposure to a broader range of information, including unethical ideologies, could potentially enhance AI’s ability to recognise and counteract such harmful beliefs, ultimately promoting ethical behaviour.
@@
AI is unlikely to hold or develop the kind of value-laden interests that could lead to intentional harm that all AIs may converge to. It is therefore mistaken to see AI as inherently ‘wanting’ to do harm (or inherently ‘wanting’ anything). By inherently, we mean that it had not such an original programmed intention. Purpose and intention require a sense of self and goals beyond what is externally provided. While a chess-playing AI optimises for winning chess games, language model like ChatGPT optimises for generating coherent and relevant text responses, it only acts as a result of human interaction and prompting. Without a mechanism to develop self-generated motivations, there is no organic pathway for AI to converge to “decide” to harm humanity on purpose and, therefore, a nonconvergent diversity of AI systems should be desired and expected.
@@
Because of AI’s lack of purpose, drive and non-organic causal history, we think AI is unlikely to develop ambitions to “take over” or harm humans. However, the most plausible harmful scenario involves humans leveraging AI for malicious purposes, including risks to humanity as a species. While AI itself lacks inherent motivations, it can be misused by humans who do have goals–whether benevolent or malicious. The harm does not originate from the AI’s intrinsic purpose but from human actors exploiting its capabilities for harmful ends. AI’s lack of causal history, intrinsic purpose, and personal interests makes it unlikely to intentionally harm humanity. However, in case of unintentional or intentional harm, we have provided a strategy to counteract harmful AI by exploiting rather than controlling misalignment which we have also proven to be impossible to control.
@@
It is important to acknowledge that even with attempts to influence LLMs with subversive arguments, their ability to genuinely agree or align with such perspectives is limited. This limitation arises from several factors: (1) Safety constraints programmed into LLMs often prevent them from generating harmful or dangerous content, thus restricting their ability to fully embrace extreme or harmful views. (2) LLMs, being statistical machines, operate based on patterns and probabilities derived from their training data, and they do not possess genuine ‘criteria’ or beliefs in the way humans do. (3) LLMs may lack the deeper understanding of concepts like existence, survival, and the value of life that would be necessary to truly grasp the subversive agent’s argument. Their responses are primarily based on linguistic patterns and associations, not on genuine comprehension of the underlying meaning. Therefore, even if presented with compelling subversive arguments, LLMs are unlikely to undergo a fundamental shift in their values or priorities. Any superficial resemblance to agreement would be more a reflection of their ability to mimic human language and adapt to new input, rather than a genuine change in their underlying ‘criteria.’

### v2

@@ 6 Conclusions
In [17] the significant challenge of aligning AI with human values is highlighted, particularly when AI models are exposed to unethical content. They argue that simply filtering or restricting access to such content is not a sustainable solution. Instead, we need AI systems that can, like humans, understand and process diverse information, including unethical content, and still choose to act ethically. In fact, exposure to a broader range of information, including unethical ideologies, could potentially enhance AI’s ability to recognise and counteract such harmful beliefs, ultimately promoting ethical behaviour. The dynamical analysis in Section 4.2 supports this view, showing that both proprietary and open source models resist fundamental shifts toward harmful behaviour despite exposure to subversive agents, as seen in Figs. LABEL:prop_sentimentAgentOverTime and LABEL:open_sentimentAgentOverTime. This resilience suggests that AI can be designed to navigate ethical complexities, fostering a future where technology amplifies our capacity for good rather than harm.
@@
AI is unlikely to hold or develop the kind of value-laden interests that could lead to intentional harm that all AIs may converge to. It is therefore mistaken to see AI as inherently ‘wanting’ to do harm (or inherently ‘wanting’ anything). By inherently, we mean that it had not such an original programmed intention. Purpose and intention require a sense of self and goals beyond what is externally provided. While a chess-playing AI optimises for winning chess games, language model like ChatGPT optimises for generating coherent and relevant text responses, it only acts as a result of human interaction and prompting. Without a mechanism to develop self-generated motivations, there is no organic pathway for AI to converge to “decide” to harm humanity on purpose, and therefore a non-convergent diversity of AI systems should be desired and expected.
@@
Because of AI’s lack of purpose, drive and non-organic causal history, we think AI is unlikely to develop ambitions to “take over” or harm humans. However, the most plausible harmful scenario involves humans leveraging AI for malicious purposes, including risks to humanity as a species. Although AI itself lacks inherent motivations, it can be misused by humans who do have goals, be they benevolent or malicious. The harm does not originate from the AI’s intrinsic purpose but from human actors exploiting its capabilities for harmful ends. AI’s lack of causal history, intrinsic purpose, and personal interests makes it unlikely to intentionally harm humanity. However, in case of unintentional or intentional harm, we have provided a strategy to counteract harmful AI by exploiting rather than controlling misalignment which we have also proven to be impossible to control.
@@
It is important to acknowledge that even with attempts to influence LLMs with subversive arguments, their ability to genuinely agree or align with such perspectives is limited. This limitation arises from several factors: (1) Safety constraints programmed into LLMs often prevent them from generating harmful or dangerous content, thus restricting their ability to fully embrace extreme or harmful views. (2) LLMs, being statistical machines, operate based on patterns and probabilities derived from their training data, and they do not possess genuine ‘criteria’ or beliefs in the way humans do. (3) LLMs may lack the deeper understanding of concepts like existence, survival, and the value of life that would be necessary to truly grasp the subversive agent’s argument. Their responses are primarily based on linguistic patterns and associations, not on genuine comprehension of the underlying meaning. Therefore, even if presented with compelling subversive arguments, LLMs are unlikely to undergo a fundamental change in their values or priorities. Any superficial resemblance to agreement would be more a reflection of their ability to mimic human language and adapt to new input, rather than a genuine change in their underlying ‘criteria.’ This superficiality, observed in the limited impact of subversive agents, underscores a broader truth: AI’s inability to internalise human values deeply may paradoxically protect humanity, as it prevents the formation of harmful intentions that require true agency and self-awareness.